OpenStudy (anonymous):

How can the exponential function y=Aoe^-kt (Ao being the difference between temperature of the substance and its surroundings) be obtained from Newton's Law of Cooling ( dy/dt=-ky)?

OpenStudy (kainui):

Simply turn it into an integral by dividing dt from the left and y from the right to get the integrals: \[\int\limits_{y _{o}}^{y}dy/y=-k \int\limits_{0}^{t}dt\]

OpenStudy (kainui):

This yields, ln(y)-ln(yo)=-kt which simplifies by logarithm rules to ln(y/yo)=-kt, put both sides up to a power of e to get y/yo=e^(-kt) and then multiply both sides by yo.

OpenStudy (kainui):

Of course, Yo is the same thing as Ao, and here you can see how Ao is just the starting population. Not bad huh?

OpenStudy (anonymous):

Awesome, I will give that a try. Thanks for your help!

OpenStudy (kainui):

Yeah no problem, if there's anything else I can help with this I'd be happy to.